OpenStudy (anonymous):

2/3=x/6=9/y To find the value of y in this proportion you must divide 54 by 2 3 4 6

OpenStudy (calculusxy):

all of them are all equivalent. so we can then see that 3, 6, 9 are all filled out as the denominator and numerator . can you find any types of pattern between them? \[\frac{ 2 }{ 3 } = \frac{ x }{ 6 } = \frac{ 9 }{ y }\]

OpenStudy (calculusxy):

hint*** look at 3 and 6..